ID |
Date |
Author |
Subject |
341
|
Tue Mar 1 05:23:02 2022 |
JM, EM | Current status |
06:20
Servo motors moved to fully out position as some work on the beam was neccessary. Beam injection had dropped from ~9e7 to around 1e7. Potentiometer reads 19.10 mm.
Still trying to improve target density. Current setpoint 100 K.
Merger OK
Data rate = 1600 kb/sec
Leakage current = 3.21 uA
System checks OK - wr fails
Base Current Difference
aida01 fault 0xab60 : 0xab8e : 46
aida02 fault 0x88d3 : 0x8900 : 45
aida03 fault 0x23aa : 0x23d8 : 46
aida04 fault 0x15d7 : 0x1605 : 46
aida05 fault 0x2f03 : 0x2f0f : 12
White Rabbit error counter test result: Passed 0, Failed 5
Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR
Avg ring pressure = 4.3E-10 mbar.
06:40
Target density still very low. Setpoint back to 116 K. Detectors still fully out, very low rutherford rate.
Attempted to move nozzle to increase density, already on maximum point.
06:50
Setpoint to 95 k. Small increase in target density, however this was temporary
|
Attachment 1: 2022-03-01_06-42-53-ExRate.png
|
|
340
|
Tue Mar 1 04:07:31 2022 |
JM, EM | Temperature cycle |
Have been unable to achieve decent target density with temperature cycling so far. Will attempt to go up to 300 k and back down to around 115 k which proved successful in previous shifts. |
339
|
Tue Mar 1 03:34:28 2022 |
JM, EM | Current status |
Temperature setpoint = 116 k
Detector IN potentiometer = 34.89 mm
Rutherford rate is reduced from previous entries. Avergae peak height appears to be ~50 counts. Fluctuations in peak height and E4 pressure are quite high. Temperature has been cycled, nozzle position has been optimised and different temperatures have been trialled however density cannot be maintained at a decent level.
Merger OK
Data rate ~ 1500 kb/sec
Leakage current = 3.2295 uA
Temperatures OK
System checks OK - wr fails
Base Current Difference
aida01 fault 0xab60 : 0xab8c : 44
aida02 fault 0x88d3 : 0x88fe : 43
aida03 fault 0x23aa : 0x23d6 : 44
aida04 fault 0x15d7 : 0x1603 : 44
aida05 fault 0x2f03 : 0x2f0f : 12
White Rabbit error counter test result: Passed 0, Failed 5
Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR
Avg ring pressure = 4.4 E-10 mbar
|
Attachment 1: 2022-03-01_04-37-03-Ex-rate.png
|
|
Attachment 2: 2022-03-01_04-41-46-correlation8.png
|
|
Attachment 3: 2022-03-01_04-41-15-goodEvents.png
|
|
Attachment 4: 2022-03-01_04-40-49-ADCitems.png
|
|
Attachment 5: 2022-03-01_04-43-53-XYtotal.png
|
|
Attachment 6: 2022-03-01_04-44-33-EyTotal.png
|
|
Attachment 7: 2022-03-01_04-45-07-ExTotal.png
|
|
338
|
Tue Mar 1 03:19:50 2022 |
JM, EM | Temperature Cycle |
04:20
Did another temperature cycle (118 K -> 125 K -> 95 K -> 116 K) trying to improve the pressures at the target and the detector count rate.
Rates are still mostly below 100 counts (Ex rate) and the pressures are still fluctuating by about an order of magnitude (E4: between 1e-9 and 1e-8 mbar). |
337
|
Tue Mar 1 02:15:29 2022 |
JM, EM | Current status |
03:15
The nozzle motor control at the target crashed, but operation has been restored.
Temperature setpoint = 116 K
Detector position IN potentionmeter = 34.89mm
Rutherford rate fluctuates, around 100 counts peak height
Merger OK
Data rate = 1356 kb/sec
Leakage current = 3.234 uA
Temperatures OK
System checks OK - wr fails
Base Current Difference
aida01 fault 0xab60 : 0xab8c : 44
aida02 fault 0x88d3 : 0x88fe : 43
aida03 fault 0x23aa : 0x23d6 : 44
aida04 fault 0x15d7 : 0x1603 : 44
aida05 fault 0x2f03 : 0x2f0f : 12
White Rabbit error counter test result: Passed 0, Failed 5
Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR
Avg ring pressure = 4.4 E-10 mbar |
Attachment 1: 2022-03-01_03-21-28_Exrate.png
|
|
Attachment 2: 2022-03-01_03-22-30_ExTotal.png
|
|
Attachment 3: 2022-03-01_03-23-06_EyTotal.png
|
|
Attachment 4: 2022-03-01_03-25-54_GoodEvents.png
|
|
Attachment 5: 2022-03-01_03-26-48_Correlations.png
|
|
Attachment 6: 2022-03-01_03-27-23_ADCdataitems.png
|
|
336
|
Mon Feb 28 23:54:02 2022 |
JM, EM | Current Status, Temperature cycle |
00:51
Temperature setpoint = 116 K
Detector position IN potentionmeter = 34.89mm
Rutherford rate fluctuates, consistently above 100 counts peak height
Merger OK
Data rate = 1600 kb/sec
Leakage current = 3.235 uA
Temperatures OK
System checks OK - wr fails
Base Current Difference
aida01 fault 0xab60 : 0xab8c : 44
aida02 fault 0x88d3 : 0x88fe : 43
aida03 fault 0x23aa : 0x23d6 : 44
aida04 fault 0x15d7 : 0x1603 : 44
aida05 fault 0x2f03 : 0x2f0f : 12
White Rabbit error counter test result: Passed 0, Failed 5
Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR
Avg ring pressure = 4.4 E-10 mbar
01:18
Rutherford rate showed sustained drop to below 10 counts peak level. E4 was also consistently down, fluctuating in the low -9 mbar range compared to -8 mbar range previously.
Temperature increased to 125 K
01:39
Completed increase to 125k, drop to 95k and back to 116k. Rate has improved somewhat, however is still lower than before. Will monitor and see if any further change in temperature helps.
01:48
E4 pressures and rutherford rate remain low ~40 counts peak height. Decreasing temperature in 1k steps to see if any improvement.
02:05
Rates still low at decreased temperature. Setting to 140 K for 15-20 mins and then bringing back down.
02:36
Temperature brought back down, optimum temperature appears to be 118 K. E4 ~ 3E-9 mbar and rutherford rate is back around 100 counts peak heught. |
Attachment 1: 2022-03-01_00-57-44-Exrate.png
|
|
Attachment 2: 2022-03-01_00-58-41-ExTotal.png
|
|
Attachment 3: 2022-03-01_00-58-14-EyTotal.png
|
|
335
|
Mon Feb 28 22:34:51 2022 |
JM, EM | Current status |
23:35
Temperature setpoint = 116 K
Detector position IN, potentiometer reads 34.89 mm
Rutherford rate fluctuates however is consistently aboe 100 counts peak height
Merger OK
Data rate = 1600 kb/sec
Leakage current = 3.23 uA
Temperatures OK
System checks OK - wr attached
Base Current Difference
aida01 fault 0xab60 : 0xab8c : 44
aida02 fault 0x88d3 : 0x88fe : 43
aida03 fault 0x23aa : 0x23d6 : 44
aida04 fault 0x15d7 : 0x1603 : 44
aida05 fault 0x2f03 : 0x2f0f : 12
White Rabbit error counter test result: Passed 0, Failed 5
Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR
|
Attachment 1: 2022-02-28_23-40-14-xyTotal.png
|
|
Attachment 2: 2022-02-28_23-39-56-Exrate.png
|
|
Attachment 3: 2022-02-28_23-39-26-goodEvents.png
|
|
Attachment 4: 2022-02-28_23-38-58-correlation8.png
|
|
Attachment 5: 2022-02-28_23-38-31-ADCitems.png
|
|
334
|
Mon Feb 28 14:47:52 2022 |
CB, TD | Monday 28 February 15:00-23:00 |
15.51 1.8.W spectra 20ms, 2ms, 200us, 20us FSR - attachments 1-8
15.55 Lost control of motors - moving DSSSDs out
15.56 new run number 79
ignore run number 80
17:58 new run number 81
CB sets motion cycle from 20.28mm to 64.36mm
Observe a second, higher energy peak during the part of the cycle where the DSSSD is moved out - see attachments 9-10
Peak appears only when moving out. Does not appear when moving in/out manually on beam storage plateau. Does not appear when moving in synched with ring, waiting less than the storage time, and moving out.
All seems to indicate it is due to beam injection and not moving out fast enough. Very weak halo piling up? Should be clear in the data depending on rate / strips affected.
18.45 new run number 82
DSSSD position fixed 34.89mm
22.10 system wide checks OK *except* WR decoder status
Base Current Difference
aida01 fault 0xab60 : 0xab8c : 44
aida02 fault 0x88d3 : 0x88fe : 43
aida03 fault 0x23aa : 0x23d6 : 44
aida04 fault 0x15d7 : 0x1603 : 44
aida05 fault 0x2f03 : 0x2f0f : 12
White Rabbit error counter test result: Passed 0, Failed 5
Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR
FEE64 temperatures OK - attachment 11
ADC data item stats OK - attachment 12
merger, tape server and console OK - no merger error messages since startup - attachments 13-15
DSSSD bias -100V leakage current-3.2360uA
run82 - integrated x & y strip energy spectra - attachments 16-17 |
Attachment 1: Screenshot_from_2022-02-28_15-51-42.png
|
|
Attachment 2: Screenshot_from_2022-02-28_15-51-27.png
|
|
Attachment 3: Screenshot_from_2022-02-28_15-50-56.png
|
|
Attachment 4: Screenshot_from_2022-02-28_15-50-44.png
|
|
Attachment 5: Screenshot_from_2022-02-28_15-50-26.png
|
|
Attachment 6: Screenshot_from_2022-02-28_15-50-08.png
|
|
Attachment 7: Screenshot_from_2022-02-28_15-49-44.png
|
|
Attachment 8: Screenshot_from_2022-02-28_15-49-28.png
|
|
Attachment 9: Screenshot_from_2022-02-28_18-03-13.png
|
|
Attachment 10: Screenshot_from_2022-02-28_18-03-20.png
|
|
Attachment 11: Screenshot_from_2022-02-28_22-13-42.png
|
|
Attachment 12: Screenshot_from_2022-02-28_22-14-12.png
|
|
Attachment 13: Screenshot_from_2022-02-28_22-15-31.png
|
|
Attachment 14: Screenshot_from_2022-02-28_22-15-22.png
|
|
Attachment 15: Screenshot_from_2022-02-28_22-15-16.png
|
|
Attachment 16: Screenshot_from_2022-02-28_22-18-23.png
|
|
Attachment 17: Screenshot_from_2022-02-28_22-18-01.png
|
|
333
|
Mon Feb 28 14:42:48 2022 |
CB, TD | Cryostat set to 115K |
Cryostat set to 115K |
332
|
Mon Feb 28 14:38:15 2022 |
TD | NewMerger statistics |
Screenshots of all NewMerger statistics |
Attachment 1: Screenshot_from_2022-02-28_15-38-04.png
|
|
Attachment 2: Screenshot_from_2022-02-28_15-37-57.png
|
|
Attachment 3: Screenshot_from_2022-02-28_15-37-52.png
|
|
Attachment 4: Screenshot_from_2022-02-28_15-37-44.png
|
|
Attachment 5: Screenshot_from_2022-02-28_15-37-32.png
|
|
Attachment 6: Screenshot_from_2022-02-28_15-37-24.png
|
|
Attachment 7: Screenshot_from_2022-02-28_15-37-10.png
|
|
Attachment 8: Screenshot_from_2022-02-28_15-37-03.png
|
|
Attachment 9: Screenshot_from_2022-02-28_15-36-55.png
|
|
Attachment 10: Screenshot_from_2022-02-28_15-36-48.png
|
|
Attachment 11: Screenshot_from_2022-02-28_15-36-39.png
|
|
Attachment 12: Screenshot_from_2022-02-28_15-36-31.png
|
|
Attachment 13: Screenshot_from_2022-02-28_15-36-23.png
|
|
Attachment 14: Screenshot_from_2022-02-28_15-36-16.png
|
|
Attachment 15: Screenshot_from_2022-02-28_15-36-07.png
|
|
Attachment 16: Screenshot_from_2022-02-28_15-35-59.png
|
|
Attachment 17: Screenshot_from_2022-02-28_15-35-51.png
|
|
Attachment 18: Screenshot_from_2022-02-28_15-35-42.png
|
|
Attachment 19: Screenshot_from_2022-02-28_15-35-36.png
|
|
Attachment 20: Screenshot_from_2022-02-28_15-35-26.png
|
|
Attachment 21: Screenshot_from_2022-02-28_15-35-14.png
|
|
Attachment 22: Screenshot_from_2022-02-28_15-35-09.png
|
|
Attachment 23: Screenshot_from_2022-02-28_15-34-52.png
|
|
Attachment 24: Screenshot_from_2022-02-28_15-34-47.png
|
|
Attachment 25: Screenshot_from_2022-02-28_15-34-42.png
|
|
Attachment 26: Screenshot_from_2022-02-28_15-34-35.png
|
|
Attachment 27: Screenshot_from_2022-02-28_15-34-28.png
|
|
Attachment 28: Screenshot_from_2022-02-28_15-34-20.png
|
|
Attachment 29: Screenshot_from_2022-02-28_15-34-12.png
|
|
Attachment 30: Screenshot_from_2022-02-28_15-34-03.png
|
|
Attachment 31: Screenshot_from_2022-02-28_15-33-50.png
|
|
Attachment 32: Screenshot_from_2022-02-28_15-33-45.png
|
|
Attachment 33: Screenshot_from_2022-02-28_15-33-39.png
|
|
Attachment 34: Screenshot_from_2022-02-28_15-33-31.png
|
|
Attachment 35: Screenshot_from_2022-02-28_15-33-24.png
|
|
Attachment 36: Screenshot_from_2022-02-28_15-33-15.png
|
|
Attachment 37: Screenshot_from_2022-02-28_15-33-07.png
|
|
331
|
Mon Feb 28 13:56:02 2022 |
JG, CB | Temperature cycle to 125K |
14:45
Rutherford rate was decreased significantly for longer period.
Temperature increased to 125 K
No large increase in chamber pressure observed.
14:55
Temperature setpoint back to 110 K.
Rutherford rate has increased. Peak height 3*10^1. |
330
|
Mon Feb 28 13:16:39 2022 |
JG, AK, CB | Current status |
14:16
beam intensity at injection: 7e7 ions
Temperature setpoint = 110 k
Rutherford rates shaky, peak max at ~ 1x10^2
Merger OK
Data rate ~ 1500 kb/sec
Ring avg pressure ~ 2.3E-10 mbar
Temperatures OK
System checks OK (wr fail)
Leakage current ~ 3.23 uA |
Attachment 1: Screenshot_from_2022-02-28_14-16-00.png
|
|
Attachment 2: Screenshot_from_2022-02-28_14-15-43.png
|
|
Attachment 3: Screenshot_from_2022-02-28_14-15-30.png
|
|
Attachment 4: Screenshot_from_2022-02-28_14-15-10.png
|
|
Attachment 5: Screenshot_from_2022-02-28_14-14-29.png
|
|
Attachment 6: Screenshot_from_2022-02-28_14-14-22.png
|
|
Attachment 7: Screenshot_from_2022-02-28_14-14-14.png
|
|
Attachment 8: Screenshot_from_2022-02-28_14-13-24.png
|
|
329
|
Mon Feb 28 11:57:22 2022 |
JG, AK, CB | Current status |
12:55
beam intensity at injection: 9e7 ions
Temperature setpoint = 110 k
Rutherford rates shaky, peak max at ~ 3x10^2
Merger OK
Data rate ~ 1400 kb/sec
Ring avg pressure ~ 2.8E-10 mbar
Temperatures OK
System checks OK (wr fail)
Leakage current ~ 3.22 uA |
Attachment 1: Screenshot_from_2022-02-28_12-56-50.png
|
|
Attachment 2: Screenshot_from_2022-02-28_12-56-40.png
|
|
Attachment 3: Screenshot_from_2022-02-28_12-56-29.png
|
|
Attachment 4: Screenshot_from_2022-02-28_12-56-16.png
|
|
Attachment 5: Screenshot_from_2022-02-28_12-55-36.png
|
|
Attachment 6: Screenshot_from_2022-02-28_12-55-08.png
|
|
Attachment 7: Screenshot_from_2022-02-28_12-54-54.png
|
|
Attachment 8: Screenshot_from_2022-02-28_12-54-33.png
|
|
328
|
Mon Feb 28 10:55:48 2022 |
JG, US | Current status |
11:55
beam intensity at injection: 8e7 ions
Temperature setpoint = 110 k
Rutherford rates shaky, peak max at ~ 3x10^2
Merger OK
Data rate ~ 1600 kb/sec
Ring avg pressure ~ 2.7E-10 mbar
Temperatures OK
System checks OK - wr fail attached
Leakage current ~ 3.22 uA |
Attachment 1: Screenshot_from_2022-02-28_11-57-17.png
|
|
Attachment 2: Screenshot_from_2022-02-28_11-57-04.png
|
|
Attachment 3: Screenshot_from_2022-02-28_11-56-48.png
|
|
Attachment 4: Screenshot_from_2022-02-28_11-54-34.png
|
|
Attachment 5: Screenshot_from_2022-02-28_11-53-25.png
|
|
Attachment 6: Screenshot_from_2022-02-28_11-52-21.png
|
|
Attachment 7: Screenshot_from_2022-02-28_11-52-14.png
|
|
Attachment 8: Screenshot_from_2022-02-28_11-51-06.png
|
|
327
|
Mon Feb 28 09:31:48 2022 |
JG, AK, CB | Current status |
10:35
beam intensity at injection: 1e8 ions
Temperature setpoint = 110 k
Rutherford rates shaky, peak max at ~ 3x10^2
Merger OK
Data rate ~ 1200 kb/sec
Ring avg pressure ~ 2.7E-10 mbar
Temperatures OK
System checks OK - wr fail attached
Leakage current ~ 3.22 uA |
Attachment 1: Screenshot_from_2022-02-28_10-31-27.png
|
|
Attachment 2: Screenshot_from_2022-02-28_10-31-16.png
|
|
Attachment 3: Screenshot_from_2022-02-28_10-31-06.png
|
|
Attachment 4: Screenshot_from_2022-02-28_10-30-49.png
|
|
Attachment 5: Screenshot_from_2022-02-28_10-30-30.png
|
|
Attachment 6: Screenshot_from_2022-02-28_10-30-24.png
|
|
Attachment 7: Screenshot_from_2022-02-28_10-30-00.png
|
|
Attachment 8: Screenshot_from_2022-02-28_10-29-26.png
|
|
326
|
Mon Feb 28 08:36:35 2022 |
JG, AK, CB | Current status |
09:35
beam intensity at injection: 8.2e7 ions
Temperature setpoint = 110 k
Rutherford rates shaky, peak max at ~ 3x10^2
Merger OK
Data rate ~ 1400 kb/sec
Ring avg pressure ~ 2.7E-10 mbar
Temperatures OK
System checks OK - wr fail attached
Leakage current ~ 3.22 uA |
Attachment 1: Screenshot_from_2022-02-28_09-36-05.png
|
|
Attachment 2: Screenshot_from_2022-02-28_09-35-56.png
|
|
Attachment 3: Screenshot_from_2022-02-28_09-35-35.png
|
|
Attachment 4: Screenshot_from_2022-02-28_09-35-18.png
|
|
Attachment 5: Screenshot_from_2022-02-28_09-35-07.png
|
|
Attachment 6: Screenshot_from_2022-02-28_09-34-21.png
|
|
Attachment 7: Screenshot_from_2022-02-28_09-34-06.png
|
|
Attachment 8: Screenshot_from_2022-02-28_09-33-48.png
|
|
325
|
Mon Feb 28 07:35:51 2022 |
JG | current logger crashed |
The current logger did crash in the evening of 25.02. and stopped logging since then, it seems.
I restarted it now (28.02. 8:30).
Seems to run stable for now.
|
Attachment 1: Screenshot_from_2022-02-28_08-38-35.png
|
|
324
|
Mon Feb 28 07:31:25 2022 |
JG, AK | Current status |
08:30
Temperature setpoint = 110 k
Rutherford rates shaky, peak max at ~ 3x10^2
Merger OK
Data rate ~ 1500 kb/sec
Ring avg pressure ~ 2.7E-10 mbar
Temperatures OK
System checks OK - wr fail attached
Leakage current ~ 3.22 uA |
Attachment 1: Screenshot_from_2022-02-28_08-30-34.png
|
|
Attachment 2: Screenshot_from_2022-02-28_08-30-13.png
|
|
Attachment 3: Screenshot_from_2022-02-28_08-29-59.png
|
|
Attachment 4: Screenshot_from_2022-02-28_08-29-44.png
|
|
Attachment 5: Screenshot_from_2022-02-28_08-29-28.png
|
|
Attachment 6: Screenshot_from_2022-02-28_08-29-17.png
|
|
Attachment 7: Screenshot_from_2022-02-28_08-29-10.png
|
|
Attachment 8: Screenshot_from_2022-02-28_08-29-04.png
|
|
323
|
Mon Feb 28 06:12:58 2022 |
JG, AK, US | Current status |
07:20
Temperature setpoint = 110 k
Rutherford rates shaky, peak max at ~ 3x10^2
Merger OK
Data rate ~ 1200 kb/sec
Ring avg pressure ~ 2.5E-10 mbar
Temperatures OK
System checks OK - wr fail attached
Leakage current ~ 3.222 uA |
Attachment 1: Screenshot_from_2022-02-28_07-16-18.png
|
|
Attachment 2: Screenshot_from_2022-02-28_07-16-02.png
|
|
Attachment 3: Screenshot_from_2022-02-28_07-15-49.png
|
|
Attachment 4: Screenshot_from_2022-02-28_07-15-28.png
|
|
Attachment 5: Screenshot_from_2022-02-28_07-12-06.png
|
|
Attachment 6: Screenshot_from_2022-02-28_07-11-59.png
|
|
Attachment 7: Screenshot_from_2022-02-28_07-11-48.png
|
|
Attachment 8: Screenshot_from_2022-02-28_07-17-27.png
|
|
322
|
Mon Feb 28 05:04:47 2022 |
JM, OF | Temperature cycle to 125K |
06:01
Rutherford rate was decreased compared to previous values. Max rate ~ 50 counts for the peak, average ~20 counts.
Temperature increased to 125 K
No large increase in chamber pressure observed.
06:25
Temperature setpoint back to 110 K.
Rutherford rate has increased, however is quite variable. Peak height 3*10^1 to 10^2 range.
06:54
Setpoint = 110 K
Ruthrford rate has stabilised around 7*10^1 to 10^2 peak height.
Merger OK
Data rate ~ 1300 kb/sec
Leakage current ~ 3.22 uA
Temperatures OK
System checks OK |
Attachment 1: 2022-02-28_06-30-39-Exrate.png
|
|